OpenStudy (anonymous):

Help me, determine whether each ordered pair is a solution of the given equation. Medal will be given. -6x+5y=-6;(1,0),(2,6/5)

OpenStudy (anonymous):

Are (1,0) and (2,6/5) the choices? First make the equation into slope intercept form. -6x + 5y = -6 5y = -6 + 6x y = 6/5x - 6/5 Does that make sense?

OpenStudy (anonymous):

Yes those are the choices

OpenStudy (anonymous):

We already know that y is 6/5 because it is the intercept. The answer is the second choice. Does that make sense?

OpenStudy (anonymous):

So the answer is 1,0 ?

OpenStudy (anonymous):

Why do you say that?

OpenStudy (anonymous):

The answer is the other choice.

OpenStudy (anonymous):

What is the other choice?

OpenStudy (anonymous):

2, 6/5